# KVQuant: Towards 10 Million Context Length LLM Inference with KV Cache Quantization

### v3

@@ Abstract.
LLMs are seeing growing use for applications which require large context windows, and with these large context windows KV cache activations surface as the dominant contributor to memory consumption during inference. Quantization is a promising approach for compressing KV cache activations; however, existing solutions fail to represent activations accurately in sub-4-bit precision. Our work, KVQuant, facilitates low precision KV cache quantization by incorporating several novel methods: (i) Per-Channel Key Quantization, where we adjust the dimension along which we quantize the Key activations to better match the distribution; (ii) Pre-RoPE Key Quantization, where we quantize Key activations before the rotary positional embedding to mitigate its impact on quantization; (iii) Non-Uniform KV Cache Quantization, where we derive per-layer sensitivity-weighted non-uniform datatypes that better represent the distributions; (iv) Per-Vector Dense-and-Sparse Quantization, where we isolate outliers separately for each vector to minimize skews in quantization ranges; and (v) Q-Norm, where we normalize quantization centroids in order to mitigate distribution shift, providing additional benefits for 2-bit quantization. By applying our method to the LLaMA, LLaMA-2, and Mistral models, we achieve $<0.1$ perplexity degradation with 3-bit quantization on both Wikitext-2 and C4, outperforming existing approaches. Our method enables serving LLaMA-7B with a context length of up to 1 million on a single A100-80GB GPU and up to 10 million on an 8-GPU system. We develop custom CUDA kernels for KVQuant, showing that we can achieve up to $\sim$1.4$\times$ speedups, compared to baseline fp16 matrix-vector multiplications, for the LLaMA-7B model. The code is available at https://github.com/SqueezeAILab/KVQuant/.
@@ 1. Introduction
Large language models (LLMs) have revolutionized many natural language processing (NLP) tasks. In order to improve the capabilities of LLMs, there is significant interest in increasing the context lengths of LLMs. Longer context lengths enable new applications, including long document summarization, retrieval for answering questions about long documents, extended multi-turn applications (chen2023longlora, ), and code analysis. To support this pull from applications, there have been significant recent advances in long-context length models in industry (anthropic, ; openai, ), as well as in academia (chen2023longlora, ).
Given the importance of LLM workloads, there is strong motivation to improve their inference efficiency. LLM inference with large context lengths can be incredibly resource-intensive; serving LLMs requires high-end GPUs, and the largest LLMs require costly multi-GPU inference setups. When analyzing the computational nature of generative inference with LLMs, it becomes quickly apparent that, for relatively small batch sizes, the computation is memory bound (kim2023squeezellm, ). With the growing divergence between computational speeds and memory speeds, this problem is only going to get worse over time (gholami2020ai_and_memory_wall, ). This makes reducing the memory bottleneck preeminently important. Further analysis shows that the memory bottleneck is strongly related to context size. For short sequence lengths, the dominant contributor to memory consumption is the weight matrices, and therefore the optimal strategy is to minimize the model size in order to reduce memory consumption as well as bandwidth requirements (kim2023full, ; kim2023squeezellm, ). However, for long sequence lengths, the main bottleneck is the memory requirements for caching Key and Value (KV) activations throughout inference. In particular, the size of the KV cache can become the dominant contributor to memory footprint, even for a 32K context limit (see Table 1), making it challenging to perform long context length inference. This challenge is further exacerbated when one considers batched inference.
@@
  • We find that the Key matrices exhibit structured outliers in specific channels before applying RoPE. However, the outlier channel magnitudes become less consistent after applying RoPE, posing a distinct challenge for low precision quantization. Based on these observations, we use per-channel quantization for Keys, and we quantize Keys before RoPE is applied (see Section 3.1 and Section 3.2).
@@
  • Even with the above, we find that outlier values in cached KV activations can significantly degrade quantization resolution. Unlike for weights, it is non-trivial to extract outlier values from activations, given the dynamic nature of activations. However, we find that we can efficiently and accurately identify and compress outlier values in order to store them compactly in a separate sparse representation. We also find that per-vector outlier detection outperforms per-matrix outlier detection with no additional memory overhead. With this method, we can attain under 0.1 perplexity degradation for 3-bit KV cache quantization on both Wikitext-2 and C4 by only removing 1% of outliers, thereby facilitating accurate inference with 4.8$\times$ longer context length.
  • For ultra low-bit precision, we find that the quantized activations can deviate significantly from their corresponding fp16 values. To address this, we propose a lightweight Q-Norm layer which shifts and scales the distribution after de-quantization to match the mean and standard deviation of corresponding fp16 values. Interestingly, the Q-Norm layer can be fused with non-uniform quantization values resulting in no overhead during inference. This was particularly helpful for 2-bit quantization (see Section 3.5).
  • Previous work has shown that the model tends to use the first token as an Attention Sink, meaning that the model allocates a large score towards the initial tokens even if they are not semantically important (xiao2023efficient, ). In our work, we demonstrate that the model is also sensitive to perturbation in the initial token. By keeping only the first token in higher precision, we can attain perplexity benefits relative to quantizing the first token, without introducing significant storage overhead (see Section 3.6).
  • We implement custom CUDA kernels to perform activation quantization efficiently during inference, achieving up to $\sim$1.4$\times$ speedups for Key and Value matrix-vector multiplications for LLaMA-7B at 4-bit precision relative to the fp16 baseline (see Section 3.8 and 4.2). These results demonstrate how our methodology allows for accurate and efficient low-bit KV cache quantization.
Figure 1. Overview of the different components used in KVQuant that result in less than 0.1 perplexity degradation over the fp16 baseline when quantizing the KV cache to 3-bit precision. As shown in  Table 2, our 3-bit approach results in 4.8×4.8$\times$4.8 × reduction in cached activation memory footprint.
@@
BS # Params Model Size (GB) fp16 KV Cache Size with different seq len (GB)
16-bit $\rightarrow$ 2-bit 32K 128K 1M 10M (16-bit $\rightarrow$ 2-bit)
@@ 2.3. KV Cache Compression
There have also been several prior works on compressing the KV cache. Some of these methods aim to only store important tokens in the KV cache and to evict less important tokens, thereby maintaining low memory usage (liu2023scissorhands, ; zhang2023h, ; ge2023model, ). Other methods aim to only retrieve a subset of tokens at each step to achieve memory bandwidth savings (ribar2023sparq, ). In this work, we explore KV cache quantization as an orthogonal direction for compressing the KV cache in order to enable long context inference.
@@ 3.1. Per-Channel Key Quantization
Existing KV cache quantization approaches perform per-token quantization (meaning that the scaling factor and zero-point are shared by elements in the same token) (sheng2023flexgen, ; zhao2023atom, ). However, due to the differing average magnitudes between channels, the values within a channel are easier to quantize when grouped together than the values across different channels. As such, to better match the distributions, we investigate per-channel KV cache quantization, meaning that the scaling factor and zero-point are shared by elements in the same channel. By sharing the scaling factor and zero-point along the channel dimension, this will naturally group together values with similar magnitudes, thereby mitigating the impacts of outlier channels on other channels when quantizing to low precision. As outlined in Appendix D, we find that per-channel quantization provides significant accuracy benefits for Keys but not for Values. By leveraging per-channel quantization for Keys and per-token quantization for Values, we observe a 3.88 perplexity improvement on Wikitext-2 for 3-bit LLaMA-7B quantization. Note that this can potentially add runtime overhead since the quantization dimension is now misaligned with the reduction dimension for the Keys when performing matrix-vector multiplications. However, we find that we are able to efficiently dequantize Keys and perform the Query-Key matrix-vector multiplication without adding runtime overhead, as shown in Section 4.2. Additionally, as outlined in Section 3.7, per-channel quantization can also be challenging due to the need to recompute scaling factors as tokens are added to the Key cache. We show that we can calibrate offline for scaling factors, thereby avoiding expensive online recomputation.
Per-channel Key quantization was also explored in another concurrent work (kivi, ), which leveraged similar intuition about grouping together large magnitude values in the same channel to minimize quantization error. Their methodology requires fine-grained grouping for per-channel quantization while maintaining a residual subset of the KV cache in fp16 (until all elements for that group have been added to the KV cache). In our work, we demonstrate that by leveraging offline calibration, we can accurately perform per-channel quantization without grouping.
@@ 3.2. Pre-RoPE Key Quantization
The Query vectors computed at each iteration will have RoPE applied ($\tilde{Q}_{m}$). When caching Key vectors, we need to either cache $\tilde{K}_{n}$, or else we need to cache $K_{n}$ and apply $R^{d}_{\theta,n}$ on-the-fly during inference. The challenge with caching Key vectors after applying this rotation is that it leads to mixing pairs of channels by different amounts for different positions in the sequence, as shown in Appendix A (since it jointly rotates pairs of channels by different angles depending on the position in the sequence). The post-RoPE activation distribution is also shown in Figure 2, demonstrating how the rotation between pairs of channels leads to less consistent channel magnitudes. This makes it harder to quantize Key activation channels which would typically have consistent large-magnitude values. This motivated our investigation into whether we could perform pre-RoPE Key quantization (meaning that we quantize $K_{n}$) and then efficiently apply the positional embeddings on-the-fly after dequantization. The benefits of pre-RoPE Key quantization are highlighted in Appendix E, yielding 0.65 perplexity improvement on Wikitext-2 for 3-bit LLaMA-7B quantization. To be able to quantize Keys pre-RoPE, we develop a fused kernel to efficiently apply RoPE post-dequantization (the details of this approach will be discussed in Section 3.8).
@@
Uniform quantization is suboptimal for KV cache quantization since the Query and Key activations are non-uniform. Additionally, KV cache loading is memory bandwidth bound, regardless of batch size or sequence length, meaning that the dequantization overhead introduced by non-uniform quantization methods is not problematic (since the added computation does not introduce any additional latency). It is therefore desirable to leverage non-uniform quantization methods for KV cache quantization.
In (kim2023squeezellm, ), the authors computed non-uniform quantization signposts using a sensitivity-weighted k-means approach. However, this is challenging to apply online during inference due to its computational cost, and it is also difficult to estimate sensitivity for activations online. We therefore facilitate efficient online non-uniform KV cache quantization by computing sensitivity-weighted quantization signposts offline on a calibration set prior to inference. Using the diagonal Fisher information matrix (derived in Appendix C), along with the quantization error for activation $A$, we formulate the error minimization objective as:
(2) $Q(A)^{*}\simeq\operatorname*{arg\,min}_{Q}\sum_{i=1}^{N}\mathcal{F}_{ii}\big{(}A-Q(A)\big{)}^{2}.$
In order to derive a per-matrix non-uniform datatype, we first normalize each vector to the range $[-1,1]$. We then minimize the objective in Equation 2 offline on a calibration set using a k-means solver in order to obtain the quantization signposts for the non-uniform datatype for each Key or Value layer. Appendix F compares our non-uniform quantization approach with existing uniform and non-uniform quantization baselines (dettmers2023qlora, ), demonstrating how our non-uniform approach provides 0.33 perplexity improvement on Wikitext-2 for LLaMA-7B relative to 3-bit uniform methods. Table 16 in Appendix J shows how computing the required Fisher information for the LLaMA-65B model takes only a few minutes, and how using the k-means solver takes only a few minutes per layer (with the computation for each layer being parallelizable). In Appendix P, we also demonstrate that we can derive a metric for accurate one-shot mixed-precision assignment (where different layers are assigned different bit widths) using the quantization error weighted by sensitivity information.
@@ 3.4. Per-Vector Dense-and-Sparse Quantization
Figure 5 in Appendix B shows the portion of elements falling within different percentiles of the dynamic range. For both Keys and Values, the majority of elements are contained within a small percentage of the dynamic range. This means that by leveraging dense-and-sparse quantization, as demonstrated in (kim2023squeezellm, ), in order to isolate a small percentage of numerical outliers, we can restrict the range that we need to represent, thereby allowing us to represent the remaining elements with greater precision.
@@
Additionally, when looking at the Key and Value distributions in Figure 2, different channels and tokens have different average magnitudes. Therefore, an element which counts as an outlier in one channel may not be an outlier in another channel (since that channel may have a greater average magnitude). It is therefore crucial to directly target the outlier values that skew the dynamic range at the granularity that we are quantizing in order to address the values that are exaggerating the range along that particular dimension. In this work, we leverage per-vector dense-and-sparse quantization, where we use a different outlier threshold per-vector (either a separate threshold per-channel for per-channel quantization, or a separate threshold per-token for per-token quantization), rather than a single outlier threshold for each layer.
@@
Note that computing outlier thresholds for per-vector dense-and-sparse quantization poses potential accuracy and efficiency challenges. However, in Section 3.7, we show that we are able to accurately calibrate for per-channel outlier thresholds offline and efficiently compute per-token outlier thresholds online. After determining the upper and lower outlier thresholds, the remaining numbers in the vector are normalized to the range $[-1,1]$, and we then minimize Equation 2 (ignoring outliers) in order to obtain the quantization signposts for the non-uniform datatype for the remaining numbers. Appendix G will demonstrate the benefits of removing a small percentage of numerical outliers and keeping them in full precision, as well as the advantages of per-vector dense-and-sparse quantization over using a single outlier threshold for each layer. As shown in Figure 1, by removing 1% of numerical outliers using per-vector outlier thresholds, we achieve an additional 0.25 perplexity improvement on Wikitext-2 for 3-bit LLaMA-7B quantization, which is within 0.08 perplexity of the fp16 baseline.
@@
Table 2. Evaluation of our method for different models using the perplexity on Wikitext-2. KVQuant results are using pre-RoPE per-channel quantization for Keys. KV cache sizes assume a sequence length of 128K (ignoring context length limits for the models). Note that ATOM and FlexGen use 4-bit quantization with group sizes of 128 and 64 with uniform quantization, respectively, and we extend their methods to 3-bit and 2-bit quantization. We incorporate Q-Norm for 2-bit experiments, and we leverage Attention-Sink Aware quantization for all bit widths. We used post-RoPE quantization for all baseline methods since it achieves higher accuracy when quantizing Keys per-token as shown in Appendix N. Table 19 in Appendix M demonstrates a full evaluation on all LLaMA, LLaMA2, and Mistral models.
@@
Perplexity KV Cache (GB) Perplexity KV Cache (GB) Perplexity KV Cache (GB) Perplexity KV Cache (GB)
@@
KVQuant-2bit 7.15 8.0 5.81 12.5 4.84 24.4 4.38 40.0
KVQuant-2bit-1% 6.02 9.3 5.37 14.5 4.36 28.3 3.72 46.5
3.5. Mitigating Distribution Shift using Q-Norm
When pushing to extremely low bit widths like 2-bit quantization, we begin to observe accuracy degradation due to distribution shift, meaning that the post-quantization distribution has a different mean and standard deviation than the pre-quantization distribution. As shown in Appendix H, this distribution shift can lead to greater error accumulation at later layers. To combat this, we introduce Q-Normalization (shortened to Q-Norm), where we normalize the quantization centroids obtained from k-means in order to ensure the post-quantization distribution has the same mean and standard deviation as the pre-quantization distribution. Our solution is inspired by (li2023norm, ), which demonstrated that ensuring that the activation distributions post-weight quantization have similar mean and standard deviation to the activation distributions pre-weight quantization can help reduce accuracy degradation. Given mean $\mu_{1}$ and standard deviation $\sigma_{1}$ for the pre-quantization distribution and mean $\mu_{2}$ and standard deviation $\sigma_{2}$ for the post-quantization distribution (computed offline on a calibration set), we normalize the quantization centroids $C_{i}$ to $\hat{C}_{i}$ as follows:
(3) $\hat{C}_{i}=\frac{(C_{i}-\mu_{2})\sigma_{1}}{\sigma_{2}}+\mu_{1}.$
During dequantization, we use $\hat{C}_{i}$ in place of $C_{i}$ such that the mean and standard deviation of dequantized values matches the original fp16 distribution. As shown in Appendix H, Q-Norm provides significant accuracy benefits for 2-bit quantization with no added inference cost.

### v4

@@ Abstract.
LLMs are seeing growing use for applications which require large context windows, and with these large context windows KV cache activations surface as the dominant contributor to memory consumption during inference. Quantization is a promising approach for compressing KV cache activations; however, existing solutions fail to represent activations accurately in sub-4-bit precision. Our work, KVQuant, facilitates low precision KV cache quantization by incorporating several novel methods: (i) Per-Channel Key Quantization, where we adjust the dimension along which we quantize the Key activations to better match the distribution; (ii) Pre-RoPE Key Quantization, where we quantize Key activations before the rotary positional embedding to mitigate its impact on quantization; (iii) Non-Uniform KV Cache Quantization, where we derive per-layer sensitivity-weighted non-uniform datatypes that better represent the distributions; and (iv) Per-Vector Dense-and-Sparse Quantization, where we isolate outliers separately for each vector to minimize skews in quantization ranges. By applying our method to the LLaMA, Llama-2, Llama-3, and Mistral models, we achieve $<0.1$ perplexity degradation with 3-bit quantization on both Wikitext-2 and C4, outperforming existing approaches. Our method enables serving LLaMA-7B with a context length of up to 1 million on a single A100-80GB GPU and up to 10 million on an 8-GPU system. We develop custom CUDA kernels for KVQuant, showing that we can achieve up to $\sim$1.7$\times$ speedups, compared to baseline fp16 matrix-vector multiplications, for the LLaMA-7B model. The code is available at https://github.com/SqueezeAILab/KVQuant/.
@@ 1. Introduction
Large language models (LLMs) have revolutionized many natural language processing (NLP) tasks. In order to improve the capabilities of LLMs, there is significant interest in increasing the context lengths of LLMs. Longer context lengths enable new applications, including long document summarization, retrieval for answering questions about long documents, extended multi-turn applications (chen2023longlora, ), and code analysis. To support this pull from applications, there have been significant recent advances in long-context length models in industry (anthropic, ; openai, ), as well as in academia (chen2023longlora, ; liu2024world, ).
Given the importance of LLM workloads, there is strong motivation to improve their inference efficiency. LLM inference with large context lengths can be incredibly resource-intensive; serving LLMs requires high-end GPUs, and the largest LLMs require costly multi-GPU inference setups. When analyzing the computational nature of generative inference with LLMs, it becomes quickly apparent that, for relatively small batch sizes, the computation is memory bound (kim2023squeezellm, ). With the growing divergence between computational speeds and memory speeds, this problem is only going to get worse over time (gholami2020ai_and_memory_wall, ). This makes reducing the memory bottleneck preeminently important. Further analysis shows that the memory bottleneck is strongly related to context size. For short sequence lengths, the dominant contributor to memory consumption is the weight matrices, and therefore the optimal strategy is to minimize the model size in order to reduce memory consumption as well as bandwidth requirements (kim2023full, ; kim2023squeezellm, ). However, as shown in Figure 1, the main bottleneck for long sequence lengths is the memory requirements for caching Key and Value (KV) activations throughout inference. This challenge is further exacerbated when one considers batched inference.
@@
  • We find that the Keys exhibit outliers in specific channels before applying RoPE. However, the outlier channel magnitudes become less consistent after applying RoPE, posing a distinct challenge for low precision quantization. We address this by quantizing Keys per-channel before RoPE is applied (see Section 3.1 and Section 3.2).
@@
  • Even with the above, we find that outlier values in cached KV activations can significantly degrade quantization resolution. Unlike for weights, it is non-trivial to extract outlier values from activations, given the dynamic nature of activations. However, we find that we can efficiently and accurately identify and compress outlier values in order to store them compactly in a separate sparse representation. We also find that per-vector outlier detection outperforms per-matrix outlier detection with no additional memory overhead. By removing only 1% of outliers, we can attain under 0.1 perplexity degradation on both Wikitext-2 and C4 for 3-bit KV cache quantization with the LLaMA, Llama-2, Llama-3, and Mistral models, thereby facilitating accurate inference with 4.8$\times$ longer context length.
  • We implement custom CUDA kernels to perform activation quantization efficiently during inference, achieving up to $\sim$1.7$\times$ speedups for Key and Value matrix-vector multiplications for LLaMA-7B at 4-bit precision relative to the fp16 baseline (see Section 3.7 and 4.4). These results demonstrate how our methodology allows for accurate and efficient low-bit KV cache quantization.
Figure 1. Overview of the different components used in KVQuant that result in less than 0.1 perplexity degradation over the fp16 baseline when quantizing the KV cache for the LLaMA-7B model to 3-bit precision. As shown in  Table 2, our 3-bit approach results in 4.8×4.8$\times$4.8 × reduction in cached activation memory footprint.
@@
BS Model Model Size (GB) fp16 KV Cache Size with Different Seq. Len. (GB)
16 $\rightarrow$ 2-bit 32K 128K 1M 10M (16 $\rightarrow$ 2-bit)
@@ 2.3. KV Cache Compression
There have also been several prior works on compressing the KV cache. Some of these methods aim to only store important tokens in the KV cache and to evict less important tokens, thereby maintaining low memory usage (liu2023scissorhands, ; zhang2023h, ; ge2023model, ; li2024snapkv, ). Other methods aim to only retrieve a subset of tokens at each step to achieve memory bandwidth savings (ribar2023sparq, ). In this work, we explore KV cache quantization as an orthogonal direction for compressing the KV cache in order to enable long context inference.
@@ 3.1. Per-Channel Key Quantization
Existing KV cache quantization approaches perform per-token quantization (meaning that the scaling factor and zero-point are shared by elements in the same token) (sheng2023flexgen, ; zhao2023atom, ). However, due to the differing average magnitudes between channels, the values within a channel are easier to quantize when grouped together than the values across different channels. As such, to better match the distributions, we investigate per-channel KV cache quantization, meaning that the scaling factor and zero-point are shared by elements in the same channel. By sharing the scaling factor and zero-point along the channel dimension, this will naturally group together values with similar magnitudes, thereby mitigating the impacts of outlier channels on other channels when quantizing to low precision. As outlined in Appendix E, we find that per-channel quantization provides significant accuracy benefits for Keys but not for Values. By leveraging per-channel quantization for Keys and per-token quantization for Values, we observe a 3.82 perplexity improvement on Wikitext-2 for 3-bit LLaMA-7B quantization. Note that this can potentially add runtime overhead since the quantization dimension is now misaligned with the reduction dimension for the Keys when performing matrix-vector multiplications. However, we find that we are able to efficiently dequantize Keys and perform the Query-Key matrix-vector multiplication without adding runtime overhead, as shown in Section 4.4. Additionally, as outlined in Section 3.6, per-channel quantization can also be challenging due to the need to recompute scaling factors as tokens are added to the Key cache. We show that we can calibrate offline for scaling factors, thereby avoiding expensive online recomputation.
Per-channel Key quantization was also explored in another concurrent work (kivi, ), which leveraged similar intuition about grouping together large magnitude values in the same channel to minimize quantization error. Their methodology requires fine-grained grouping for per-channel quantization while maintaining a residual subset of the KV cache in fp16. In our work, we demonstrate that by leveraging offline calibration, we can accurately perform per-channel quantization without grouping.
@@ 3.2. Pre-RoPE Key Quantization
When caching Key vectors, we therefore need to either cache $\tilde{K}_{n}$, or else we need to cache $K_{n}$ and apply $R^{d}_{\theta,n}$ on-the-fly during inference. The challenge with caching Key vectors after applying this rotation is that it leads to mixing pairs of channels by different amounts for different positions in the sequence, as shown in Appendix A (since it jointly rotates pairs of channels by different angles depending on the position in the sequence). The post-RoPE activation distribution is also shown in Figure 2, demonstrating how the rotation between pairs of channels leads to less consistent channel magnitudes. This makes it harder to quantize Key activation channels which would typically have consistent large-magnitude values. This motivated our investigation into whether we could perform pre-RoPE Key quantization (meaning that we quantize $K_{n}$) and then efficiently apply the positional embeddings on-the-fly after dequantization. The benefits of pre-RoPE Key quantization are highlighted in Appendix F, yielding 0.82 perplexity improvement on Wikitext-2 for 3-bit LLaMA-7B quantization. To be able to quantize Keys pre-RoPE, we develop a fused kernel to efficiently apply RoPE post-dequantization (the details of this approach will be discussed in Section 3.7).
@@
Uniform quantization is suboptimal for KV cache quantization since the Query and Key activations are non-uniform. Additionally, KV cache loading is memory bandwidth bound, regardless of batch size or sequence length, meaning that the dequantization overhead introduced by non-uniform quantization methods is not problematic (since the added computation does not introduce any additional latency). It is therefore desirable to leverage non-uniform quantization methods for KV cache quantization. In (kim2023squeezellm, ), the authors computed non-uniform quantization signposts using a sensitivity-weighted k-means approach. However, this cannot be applied directly to KV cache quantization as the Values are quantized dynamically at runtime, which means that we would need to apply K-means online during inference, and it is also difficult to estimate sensitivity for activations online.
We therefore facilitate efficient online non-uniform KV cache quantization by deriving a per-tensor non-uniform datatype offline on calibration data, which is then rescaled per-channel or per-token to accurately represent the key and value distributions. We compute sensitivity-weighted quantization signposts offline on a calibration set prior to inference, while maintaining compatibility with per-vector quantization by separately normalizing each channel to the range $[-1,1]$ prior to deriving the shared datatype. Using the diagonal Fisher information matrix (derived in Appendix B), along with the quantization error for activation $A$, we formulate the error minimization objective as follows, where $A$ is flattened to one dimension and where $N$ is the number of elements from all of the samples in our calibration set:
(2) $Q(A)^{*}\simeq\operatorname*{arg\,min}_{Q}\sum_{i=1}^{N}\mathcal{F}_{ii}\big{(}A_{i}-Q(A_{i})\big{)}^{2}.$
We modify the objective in Equation 2 as described in Appendix C in order to apply it using the normalized activation values. We then minimize it offline on a calibration set using a k-means solver in order to obtain the quantization signposts for the non-uniform datatype for each Key or Value layer. Appendix G compares our non-uniform quantization approach with existing uniform and non-uniform quantization baselines (dettmers2023qlora, ), demonstrating how our non-uniform approach provides 0.29 perplexity improvement on Wikitext-2 for LLaMA-7B relative to 3-bit uniform methods. Table 14 in Appendix J shows how computing the required Fisher information for the LLaMA-65B model takes only a few minutes, and how using the k-means solver takes only a few minutes per layer (with the computation for each layer being parallelizable).
@@ 3.4. Per-Vector Dense-and-Sparse Quantization
As shown in Figure 5 in Appendix D, for both Keys and Values, the majority of elements are contained within a small percentage of the dynamic range. This means that by leveraging dense-and-sparse quantization, as demonstrated in (kim2023squeezellm, ), in order to isolate a small percentage of numerical outliers, we can restrict the range that we need to represent, thereby allowing us to represent the remaining elements with greater precision. However, when looking at the Key and Value distributions in Figure 2, different channels and tokens have different average magnitudes. Therefore, an element which counts as an outlier in one channel may not be an outlier in another channel (since that channel may have a greater average magnitude), making naive application of dense-and-sparse quantization suboptimal. It is therefore crucial to directly target the outlier values that skew the dynamic range at the granularity that we are quantizing in order to address the values that are exaggerating the range along that particular dimension. In this work, we leverage per-vector dense-and-sparse quantization, where we use a different outlier threshold per-vector (either a separate threshold per-channel for per-channel quantization, or a separate threshold per-token for per-token quantization), rather than a single outlier threshold for each layer.
@@
Note that computing outlier thresholds for per-vector dense-and-sparse quantization poses potential accuracy and efficiency challenges. However, in Section 3.6, we show that we are able to accurately calibrate for per-channel outlier thresholds offline and efficiently compute per-token outlier thresholds online. After determining the upper and lower outlier thresholds, the remaining numbers in the vector are normalized to the range $[-1,1]$, and we then minimize Equation 2 (ignoring outliers) in order to obtain the quantization signposts for the non-uniform datatype for the remaining numbers. Appendix H will demonstrate the benefits of removing a small percentage of numerical outliers and keeping them in full precision, as well as the advantages of per-vector dense-and-sparse quantization over using a single global outlier threshold for each layer. As shown in Figure 1, by removing 1% of numerical outliers using per-vector outlier thresholds, we achieve an additional 0.19 perplexity improvement on Wikitext-2 for 3-bit LLaMA-7B quantization, which is within 0.07 perplexity of the fp16 baseline.
@@
Table 2. Evaluation of our method for different models using the perplexity (PPL) on Wikitext-2. KVQuant results are using pre-RoPE per-channel quantization for Keys. KV cache sizes are estimated assuming a sequence length of 128K (ignoring context length limits for the models). Note that ATOM and FlexGen use 4-bit quantization with group sizes of 128 and 64 with uniform quantization, respectively, and we extend their methods to 3-bit and 2-bit quantization. We leverage Attention Sink-Aware quantization for all bit widths. We used post-RoPE quantization for all baseline methods since it achieves higher accuracy when quantizing Keys per-token as shown in Appendix N. Table 16 in Appendix M demonstrates a full evaluation on all LLaMA, Llama-2, Llama-3, and Mistral models.
@@
PPL KV Cache (GB) PPL KV Cache (GB) PPL KV Cache (GB) PPL KV Cache (GB)
@@
KVQuant-2bit 7.23 8.0 5.82 12.5 4.87 24.4 4.03 40.0
KVQuant-2bit-1% 6.01 9.3 5.36 14.5 4.35 28.3 3.70 46.5